OpenStudy (anonymous):

What is the vertex form of the equation? y=x^2+4x-3 Can you explain how to do this?

OpenStudy (anonymous):

To start off, let's factor out the two from 2x^2 + 4x. y = 2*(x^2 + 2x) - 3

OpenStudy (anonymous):

Now, we need to complete the square for what is in the parenthesis, so that we can get something off the form (x-h)^2. y = 2*(x^2 + 2x + 1) - 3 BUT, we added a 1 in the parenthesis, which when multiplied by the two adds 2 to the right hand side, so we need to add 2 to the left hand side also, giving us:

OpenStudy (anonymous):

y + 2 = 2*(x^2 + 2x + 1) - 3 y + 2 = 2*(x + 1)^2 - 3 y = 2*(x+1)^2 - 5 And there you go. :)

OpenStudy (anonymous):

thank u